OpenStudy (anonymous):

Two equations are shown: Equation 1: 3/4 (x − 12)=12 Equation 2: 3/4 y − 12=12 Solve each equation. Then, enter a number in each box to make this statement true. The value of x is: The value of y is:

OpenStudy (anonymous):

Can you help explain so I know how to it do in case I get another similar question like this?

Vocaloid (vocaloid):

the basic rule is just to get x by itself start with Equation 1: 3/4 (x − 12)=12 distribute the (3/4) (3/4)(x-12) = ?

Vocaloid (vocaloid):

uh not quite...

Vocaloid (vocaloid):

hint: A(B-C) = A*B - A*C (3/4)(x-12) = _________ follow the pattern

OpenStudy (anonymous):

9? im not sure im having a hard time understanding these questions sorry if i'm making you overwhelmed or anything its not my intentions honestly.

Vocaloid (vocaloid):



Vocaloid (vocaloid):



Vocaloid (vocaloid):

meaning that A = 3/4, B = x, C = 12 now find A*B - A*C

OpenStudy (anonymous):

so 3/4 * x and 3/4 *12

Vocaloid (vocaloid):

yes that gives us (3/4)x- 9 now we have (3/4)x - 9 = 12 now solve for x

OpenStudy (anonymous):

o:

Vocaloid (vocaloid):

(3/4)x - 9 = 12 add 9 to both sides

OpenStudy (anonymous):

to 3/4 and 9?

Vocaloid (vocaloid):

no, you can't add 9 to (3/4) since (3/4) is attached to x

Vocaloid (vocaloid):

(3/4)x - 9 + 9 = ?

OpenStudy (anonymous):

(3/4)x - 18 = 12

Vocaloid (vocaloid):

nope -9 + 9 = ?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

good. so what is (3/4)x - 9 + 9 ?

OpenStudy (anonymous):

(3/4)x = 12??

Vocaloid (vocaloid):

almost you also need to add 9 to the right side

Vocaloid (vocaloid):

(3/4)x - 9 = 12 becomes (3/4)x = 12 + 9 = 21 with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

good (3/4)x = 21 now solve for x

OpenStudy (anonymous):

28

Vocaloid (vocaloid):

good. so x = 28 ready for equation 2?

OpenStudy (anonymous):

Thank you and yes

Vocaloid (vocaloid):

3/4 y − 12=12 again, we want to get y by itself start by adding 12 to both sides

OpenStudy (anonymous):

3/4 y - 12 + 12

OpenStudy (anonymous):

so um 3/4 y - 12 = 12

OpenStudy (anonymous):

@Vocaloid am i right or am i speaking nonsense?

Vocaloid (vocaloid):

not quite (3/4)y - 12 + 12 = ??

OpenStudy (anonymous):

sorry @Vocaloid connection is bad right now but i fixed it um (3/4)y - 12??

OpenStudy (anonymous):

am i right? sorry for bothering you i just need help @Vocaloid

Vocaloid (vocaloid):

the - 12 and the 12 cancel out... (3/4)y - 12 + 12 = 12 + 12 (3/4)y = 24 solve for y

OpenStudy (anonymous):

32 @Vocaloid

Vocaloid (vocaloid):

good

OpenStudy (anonymous):

Thank you @Vocaloid for your patience and kindness (: take this medal as my appreciate

Vocaloid (vocaloid):

no problem, feel free to tag me anytime if I'm online